# Gradient Flow Finite Element Discretisations with Energy-Based $hp$–Adaptivity for the Gross-Pitaevskii Equation with Angular Momentum Rotation

### v1

@@ Abstract.
This article deals with the stationary Gross-Pitaevskii non-linear eigenvalue problem in the presence of a rotating magnetic field that is used to model macroscopic quantum effects such as Bose-Einstein condensates (BECs). In this regime, the ground-state wave-function can exhibit an a priori unknown number of quantum vortices at unknown locations, which necessitates the exploitation of adaptive numerical strategies. To this end, we consider the conforming finite element method and introduce a combination of a Sobolev gradient descent that respects the energy-topology of the problem to solve the non-linearity and an $hp$-adaptive strategy that is solely based on energy decay rather than a posteriori error estimators for the refinement process. Numerical results demonstrate that the $hp$-adaptive strategy is highly efficient in terms of accuracy to compute the ground-state wave function and energy for several test problems where we observe exponential convergence.
@@
Variational PDE, linear and non-linear eigenvalue problems, semi-linear elliptic operators, Gross-Pitaevskii equation, momentum rotation, energy minimisation, gradient flows, iterative Galerkin procedures, adaptive finite element methods, $hp$–adaptivity.
@@ 1. Introduction
In quantum physics, the Gross–Pitaevskii equation (GPE) [22, 15, 18] is widely used to study macroscopic quantum effects such as Bose–Einstein condensates (BEC), superconductivity or superfluidity. In its physical interpretation of the state of a collection of bosonic particles at zero Kelvin, or very small temperature, it can be rigorously derived from the many-body Schrödinger equation as the wave function collapses to a symmetric tensor product of a single-particle function. Such a Hartree–Fock ansatz becomes exact in the (dilute) mean-field limit, see [33, 23, 32] for some rigorous results.
The GPE exists both in its time-dependent and steady-state form. Within this work, we focus on the latter one with the presence of an external stirring magnetic field with magnetic vector potential $\bm{\mathsf{A}}$ that can trigger the formation of quantum vortices with a quantized vorticity. Since the number of such vortices, as well as their location, is unknown a priori, the use of adaptive finite element strategies is highly desirable. In this paper, we present a novel $hp$–adaptive finite element gradient flow method to minimise the Gross–Pitaevskii energy functional with the inclusion of a rotating magnetic field, given by
@@
First, early contributions considered the imaginary time method originating from the physics literature, see, e.g., [1] where the critical angular velocities that lead to vortex formation are studied. Later, contributions of Bao et al. [14, 12, 11] introduced a mathematical framework to tackle the problem directly by energy minimisation. Different metrics to consider the Sobolev gradient of the energy functional have then been introduced in [19], while [20, 7] considers conjugate gradient type methods. In [17], the authors consider a second-order flow involving the second time-derivative of the energy-functional. Very recently, Henning and Yadav contributed substantially to the analysis of the gradient descent method [29], the finite-element a priori convergence analysis [30] and the numerical study of conjugate-type gradient descents [2]. Finally, we note that the problem can also be tackled by Newton-type methods, see, e.g., [37, 3] and [38] limited to the non-rotating case.
@@
Second, the ground-state solution $u$ can be discretised in different ways covered by standard techniques such as the finite difference [11, 19] or the finite element method [1, 14, 12, 19, 2, 29, 24], as well as (pseudo-) spectral methods [11, 7, 20, 17]. In the spirit of our proposed $hp$-strategy, which aims to design highly accurate approximation spaces with few number of degrees of freedom, are the recent approaches [35, 27] based on the Localized Orthogonal Decomposition (LOD) for the case with and without rotating magnetic field, respectively.
@@
In this paper, we propose a new $hp$-adaptive finite-element strategy to solve the non-linear eigenvector problem corresponding to the critical points of the energy functional (1). Our approach relies on a combination of the metric-adaptive gradient flow [28, 29] and an energy-specific refinement strategy inspired from [25, 31] (see also the related local error reduction approach [8] for $hp$–adaptive discretisations of general linear elliptic problems). The energy-based refinement strategy is considerably simpler compared to a residual-based a posteriori criteria. Further, it is important to note that within an adaptive strategy, the gradient descent can be combined with the adaptive procedure merging both iterative processes into a single unified one. With respect to [25] we present three key novel developments. First, we extend the GPE to the case with rotating magnetic field and, second, also present new analytical results on the convergence of the gradient flow (Theorem 2.6). Third, we propose a new $hp$-adaptive strategy leading to much faster (exponential) convergence with respect to the number of degrees of freedom. Mesh-adaptive strategies are particularly suited in this case, since the number and location of the vortices is unknown a priori. Furthermore, high-order approximations enable the computation of very highly accurate ground-states and thus high-fidelity in the computed energies.
@@
The outline of this paper is as follows. In Section 2, we begin by specifying the problem setting and recalling the framework of the metric-adaptive Sobolev gradient approach. Then, we prove some novel results for the gradient flow, cf. Theorem 2.6. Section 3 presents the $hp$-adaptive strategy, while Section 4 presents numerical experiments illustrating the performance of the proposed method. Finally, in Section 5 we summarise the work presented in this article and highlight potential future areas of research.
@@ 2. Projected Sobolev gradient flow
In this section, we introduce the Gross–Pitaevskii equation for the associated minimisation problem given in (7) below, and establish the continuous and discrete projected Sobolev gradient flow for its exact and numerical solution, respectively.
@@
    where $\bm{\mathsf{x}}=(x,y)\in\mathbb{R}^{2}$ and $\bm{\mathsf{x}}=(x,y,z)\in\mathbb{R}^{3}$, respectively, denote the Euclidean coordinates, and ||$\cdot$|| ⋅ | the Euclidean norm in $\mathbb{R}^{d}$, for space dimensions $d=2,3$.
@@
$\bm{\mathsf{A}}(\bm{\mathsf{x}}):=\begin{cases}(y,-x)^{\top}&\text{for $\bm{\mathsf{x}}=(x,y)\in\Omega\subset\mathbb{R}^{2}$},\\ (y,-x,0)^{\top}&\text{for $\bm{\mathsf{x}}=(x,y,z)\in\Omega\subset\mathbb{R}^{3}$},\end{cases}$ (5)
@@
In particular, this implies that the energy functional from (1) only takes real values, meaning that
$\{\mathsf{E}(v):\,v\in\mathbb{H}\}\subset\mathbb{R}.$ (6)
It can further be shown that there exists an element in $\mathcal{S}_{\mathbb{H}}$, which is not necessarily unique, at which the Gross–Pitaevskii functional $\mathsf{E}$ from (1) restricted to $\mathcal{S}_{\mathbb{H}}$ attains its minimum; we provide a proof of this result in Appendix A, thereby generalising the existence theorems of [13, 10]. In turn, the problem
$\mathsf{E}(u_{\mathrm{GS}}):=\min_{v\in\mathcal{S}_{\mathbb{H}}}\mathsf{E}(v),$ (7)
which is the focus of this paper, is well-posed (up to uniqueness). We will refer to any global minimiser of the Gross–Pitaevskii functional on the sphere $\mathcal{S}_{\mathbb{H}}$ as a ground-state.
2.2. The Gross–Pitaevskii equation (GPE)
@@
with $\mathbb{H}^{\star}$ signifying the dual space of $\mathbb{H}$ (composed of all bounded linear forms on $\mathbb{H}$ that map into $\mathbb{R}$), and $\langle\cdot,\cdot\rangle:\,\mathbb{H}^{\star}\times\mathbb{H}\to\mathbb{R}$ the associated dual product. A simple calculation employing the Riesz representation theorem reveals that
$\langle\mathsf{E}^{\prime}(u),{v}\rangle=\Re\int_{\Omega}\left(\nabla u\cdot\nabla\overline{v}+2V(\bm{\mathsf{x}})u\overline{v}+2\beta|u|^{2}u\overline{v}-2i\omega\overline{v}(\bm{\mathsf{A}}(\bm{\mathsf{x}})\cdot\nabla u)\right)\,\mathsf{d}\bm{\mathsf{x}},$ (10)
@@
$\Re\int_{\Omega}\left(\nabla u\cdot\nabla\overline{v}+2V(\bm{\mathsf{x}})u\overline{v}+2\beta|u|^{2}u\overline{v}-2i\omega\overline{v}(\bm{\mathsf{A}}(\bm{\mathsf{x}})\cdot\nabla u)\right)\,\mathsf{d}\bm{\mathsf{x}}=\lambda\Re\left(\int_{\Omega}u\overline{v}\,\mathsf{d}\bm{\mathsf{x}}\right)$ (11)
for all $v\in\mathbb{H}$. Equivalently, see [30, Prop. A.1], a solution $(\lambda,u)$ of (11) satisfies
@@
meaning that it does not matter whether or not the real part is applied. This is the weak form of the stationary Gross-Pitaevskii equation (GPE) with angular momentum, which can also be written as a semi-linear eigenfunction problem, namely, in strong form $(\lambda,u)$ satisfy
$-\Delta u+2V(\bm{\mathsf{x}})u+2\beta|u|^{2}u-2i\omega(\bm{\mathsf{A}}(\bm{\mathsf{x}})\cdot\nabla u)=\lambda u$ (13)
in $\mathbb{H}^{\star}$. An $\mathrm{L}^{2}(\Omega)$-normalised solution $u\in\mathbb{H}$ of (13) with minimal energy is termed a ground-state, denoted by $u_{\mathrm{GS}}\equiv u$, cf. (7), whereas any $\mathrm{L}^{2}(\Omega)$-normalised $u\in\mathbb{H}$ solving (13), with an energy strictly higher than the ground-state energy, i.e., $\mathsf{E}(u)>\mathsf{E}(u_{\mathrm{GS}})=\min_{v\in\mathcal{S}_{\mathbb{H}}}\mathsf{E}(v),$ is called an excited state.
@@
2.3. Projected Sobolev gradient flow
In this section we aim to establish a projected Sobolev gradient flow formulation for the minimisation of the Gross–Pitaevskii energy functional $\mathsf{E}$ from (1). For this purpose, we will consider the real framework of the Gross–Pitaevskii equation (12) rather than the complex Hilbert space approach. We emphasise, however, that both avenues ultimately result in the identical gradient flow.
@@ 2.3.1. Weighted energy inner-product
cf. [29, Eq. (7)], which will be investigated in the ensuing result.
@@
For any given $z\in\mathbb{H}$, the map $a_{z}:\mathbb{H}\times\mathbb{H}\to\mathbb{C}$ from (14) defines a (complex) inner-product on the space $\mathbb{H}=\mathrm{H}_{0}^{1}(\Omega;\mathbb{C})$. Moreover, the induced norm is equivalent to the $\mathbb{H}$-norm defined in (3). In particular, for a constant $c>0$ independent of $z\in\mathbb{H}$, the uniform coercivity property $a_{z}(u,u)\geq c\left\|\nabla u\right\|_{\mathrm{L}^{2}(\Omega)}^{2}$ for all $u\in\mathbb{H}$ holds.
@@
We will recall several ideas from [19, §2.1]. Throughout, let $z\in\mathbb{H}$ be arbitrary.
@@
    Bilinearity and symmetry. First, we note that the mapping $u\mapsto a_{z}(u,v)$ is linear for any fixed $v\in\mathbb{H}$. Moreover, we have that $a_{z}(u,v)=\overline{a_{z}(v,u)}$ for any $u,v\in\mathbb{H}$. Indeed, since $\beta\in\mathbb{R}$ and $V\in\mathrm{L}^{\infty}(\Omega;\mathbb{R})$, the conjugate-symmetry is directly observed for the first three terms in the integral from (14). Moreover, for the last term, we exploit the fact that $u$ and $v$ vanish along the boundary of $\Omega$, and hence application of the divergence theorem implies that
    $-i\omega\int_{\Omega}\overline{v}(\bm{\mathsf{A}}\cdot\nabla u)\,\mathsf{d}\bm{\mathsf{x}}=i\omega\int_{\Omega}u\operatorname*{div}(\overline{v}\bm{\mathsf{A}})\,\mathsf{d}\bm{\mathsf{x}}=i\omega\int_{\Omega}u(\nabla\overline{v}\cdot\bm{\mathsf{A}})\,\mathsf{d}\bm{\mathsf{x}}=\overline{-i\omega\int_{\Omega}\overline{u}(\bm{\mathsf{A}}\cdot\nabla v)\,\mathsf{d}\bm{\mathsf{x}}},$
@@
    $\varepsilon:=\frac{\omega^{2}r_{\Omega}^{2}}{\omega^{2}r_{\Omega}^{2}+2\delta_{\Omega}}\in(0,1),$
    with $\delta_{\Omega}>0$ from (4), it follows that $2V(\bm{\mathsf{x}})-\varepsilon^{-1}\omega^{2}r^{2}_{\Omega}\geq 0$, for (almost) all $\bm{\mathsf{x}}\in\Omega$. Therefore, exploiting that $\beta\geq 0$, we have
@@
    $c:=1-\varepsilon=\frac{2\delta_{\Omega}}{\omega^{2}r_{\Omega}^{2}+2\delta_{\Omega}}>0;$ (19)

### v2

# Gradient Flow Finite Element Discretisations with Energy-Based hphp–Adaptivity for the Gross-Pitaevskii Equation with Angular Momentum

@@ Abstract.
This article deals with the stationary Gross-Pitaevskii non-linear eigenvalue problem in the presence of a rotating magnetic field that is used to model macroscopic quantum effects such as Bose-Einstein condensates (BECs). In this regime, the ground-state wave-function can exhibit an a priori unknown number of quantum vortices at unknown locations, which necessitates the exploitation of adaptive numerical strategies. To this end, we consider the conforming finite element method in combination with a discrete Sobolev gradient descent, which is guided by the energy-topology of the problem, to address the nonlinearity. In addition, a key novelty of this work is an $hp$-adaptive strategy that is solely based on energy decay rather than a posteriori error estimators for the refinement process. Numerical results demonstrate that the $hp$-adaptive strategy is highly efficient in terms of accuracy to compute the ground-state wave function and energy for several test problems where we observe exponential convergence.
@@
Variational PDE, linear and non-linear eigenvalue problems, semi-linear elliptic operators, Gross-Pitaevskii equation, angular momentum, energy minimisation, gradient flows, iterative Galerkin procedures, adaptive finite element methods, hphp–adaptivity.
@@ 1. Introduction
In quantum physics, the Gross–Pitaevskii equation (GPE) [22, 15, 17] is widely used to study macroscopic quantum effects such as Bose–Einstein condensates (BECs), superconductivity or superfluidity. In its physical interpretation of the state of a collection of bosonic particles at zero Kelvin, or very small temperature, it can be rigorously derived from the many-body Schrödinger equation as the wave function collapses to a symmetric tensor product of a single-particle function. Such a Hartree–Fock ansatz becomes exact in the (dilute) mean-field limit, see [34, 23, 33] for some rigorous results.
The GPE exists both in its time-dependent and steady-state form. Within this work, we focus on the latter one with the presence of an external stirring magnetic field with magnetic vector potential $\bm{\mathsf{A}}$ that can trigger the formation of quantum vortices with a quantized vorticity. Since the number of such vortices, as well as their location, is unknown a priori, the use of adaptive finite element strategies is highly desirable. In this paper, we present a novel hphp–adaptive finite element gradient flow method to minimise the Gross–Pitaevskii energy functional with the inclusion of a rotating magnetic field, given by
@@
First, early contributions considered the imaginary time method originating from the physics literature, see, e.g., [1] where the critical angular velocities that lead to vortex formation are studied. Later, contributions of Bao et al. [14, 12, 11] introduced a mathematical framework to tackle the problem directly by energy minimisation, recently further analysed in [24]. Different metrics to consider the Sobolev gradient of the energy functional have then been introduced in [18], while [19, 7] considers conjugate gradient type methods. In [16], the authors consider a second-order flow involving the second time-derivative of the energy-functional. Very recently, Henning and Yadav contributed substantially to the analysis of the gradient descent method [30], the finite-element a priori convergence analysis [31] and the numerical study of conjugate-type gradient descents [2]. Finally, we note that the problem can also be tackled by Newton-type methods, see, e.g., [37, 3] and [38] limited to the non-rotating case.
@@
Second, the ground-state solution $u$ can be discretised in different ways covered by standard techniques such as the finite difference [11, 18] or the finite element method [1, 14, 12, 18, 2, 30, 25], as well as (pseudo-) spectral methods [11, 7, 19, 16]. In the spirit of our proposed $hp$-strategy, which aims to design highly accurate approximation spaces with few number of degrees of freedom, are the recent approaches [35, 28, 21] based on the Localized Orthogonal Decomposition (LOD) for the case with and without rotating magnetic field, respectively and the related Ginzburg-Landau equations.
@@
In this paper, we propose a new $hp$-adaptive finite-element strategy to solve the non-linear eigenvector problem corresponding to the critical points of the energy functional (1). Our approach relies on a combination of the metric-adaptive gradient flow [29, 30] and an energy-specific refinement strategy inspired from [26, 32] (see also the related local error reduction approach [8] for hphp–adaptive discretisations of general linear elliptic problems). The energy-based refinement strategy is considerably simpler compared to a residual-based a posteriori criteria. Further, it is important to note that within an adaptive strategy, the gradient descent can be combined with the adaptive procedure merging both iterative processes into a single unified one. With respect to [26], we introduce three key novel developments. First, we extend the GPE to the setting of a rotating magnetic field. Second, we establish a convergence result for a discrete gradient-flow iteration (Theorem 2.7) that is derived from an energy-related continuous projected Sobolev gradient flow; these ideas follow closely the work [29], where the rotation-free case has been considered. Third, we propose a new $hp$-adaptive strategy that yields substantially faster (exponential) convergence with respect to the number of degrees of freedom.
Mesh-adaptive strategies are particularly suited in this case, since the number and location of the vortices is unknown a priori. Furthermore, high-order approximations enable the computation of highly accurate ground-states and thus high-fidelity in the computed energies.
@@
The outline of this paper is as follows. In §2, we begin by specifying the problem setting and recalling the framework of the metric-adaptive Sobolev gradient approach; we also prove some instrumental results for the discrete gradient flow, cf. Theorem 2.7. In §3 we present the $hp$-adaptive strategy, while in §4 we conduct some numerical experiments illustrating the superior performance of the proposed energy-based $hp$-adaptive approach. Finally, in §5 we summarise the work presented in this article and highlight potential future areas of research.
@@ 2. Projected Sobolev gradient flow
In this section, we introduce the GPE for the associated minimisation problem given in (7) below, and establish the continuous and discrete projected Sobolev gradient flow for its exact and numerical solution, respectively.
@@
    where $\bm{\mathsf{x}}=(x,y)\in\mathbb{R}^{2}$ and $\bm{\mathsf{x}}=(x,y,z)\in\mathbb{R}^{3}$, respectively, denote the Euclidean coordinates, and $|\cdot|$ the Euclidean norm in $\mathbb{R}^{d}$, for space dimensions $d=2,3$.
@@
$\bm{\mathsf{A}}(\bm{\mathsf{x}}):=\begin{cases}(y,-x)^{\intercal}&\text{for $\bm{\mathsf{x}}=(x,y)\in\Omega\subset\mathbb{R}^{2}$},\\ (y,-x,0)^{\intercal}&\text{for $\bm{\mathsf{x}}=(x,y,z)\in\Omega\subset\mathbb{R}^{3}$},\end{cases}$ (5)
@@
In particular, in view of (4), this implies that the energy functional from (1) only takes non-negative real values, meaning that
$\{\mathsf{E}(v):\,v\in\mathbb{H}\}\subset\mathbb{R}_{\geq 0}.$ (6)
The focus of this paper is the numerical approximation of a global minimiser of the energy functional $\mathsf{E}$ on the sphere $\mathcal{S}_{\mathbb{H}}$, which we refer to as a ground-state of the Gross-Pitaevskii energy (1). The following result generalises the existence theorems of [13, 10].
Proposition 2.1.
Under the assumption (V1) above there exists a (non-unique) $u_{\mathrm{GS}}\in\mathcal{S}_{\mathbb{H}}$, which satisfies
$\mathsf{E}(u_{\mathrm{GS}})=\min_{v\in\mathcal{S}_{\mathbb{H}}}\mathsf{E}(v)\geq 0;$ (7)
i.e., in particular, the Gross-Potaevskii energy functional $\mathsf{E}$ from (1) restricted to $\mathcal{S}_{\mathbb{H}}$ attains its minimum.
Proof.
From (6) we notice that the energy functional $\mathsf{E}$ is bounded from below. In addition, based on the condition (V1), it is coercive (to be shown later on, see (22) below). Therefore, there exists a bounded sequence $\{u_{n}\}_{n}\subset\mathcal{S}_{\mathbb{H}}$ with
$\lim_{n\to\infty}\mathsf{E}(u_{n})=\inf_{v\in\mathcal{S}_{\mathbb{H}}}\mathsf{E}(v).$
By the reflexivity of $\mathbb{H}$ and exploiting the compact embedding $\mathbb{H}\hookrightarrow\mathrm{L}^{2}(\Omega)$ we can extract a subsequence $\{u_{n_{k}}\}_{k}$ that converges weakly in $\mathbb{H}$ and strongly in $\mathrm{L}^{2}(\Omega)$ to some element $u\in\mathbb{H}$. Due to the convergence in $\mathrm{L}^{2}(\Omega)$, we indeed have that $\left\|u\right\|_{\mathrm{L}^{2}(\Omega)}=1$; i.e., $u\in\mathcal{S}_{\mathbb{H}}$. We further remark that the energy functional $\mathsf{E}$ is convex, which follows from the same arguments as in the proof of the coercivity from Proposition 2.3 below, and strongly continuous. In turn, the functional $\mathsf{E}$ is weakly sequentially lower semicontinuous; see, e.g., [39, Prop. 25.20]. Consequently, we have that
thereby implying that $u$ is indeed a minimiser of the energy functional $\mathsf{E}$ on $\mathcal{S}_{\mathbb{H}}$. ∎
2.2. The Gross–Pitaevskii equation
@@
[…]
$\langle\mathsf{E}^{\prime}(u),{v}\rangle=\Re\left(\int_{\Omega}\left(\nabla u\cdot\nabla\overline{v}+2V(\bm{\mathsf{x}})u\overline{v}+2\beta|u|^{2}u\overline{v}-2i\omega\overline{v}(\bm{\mathsf{A}}(\bm{\mathsf{x}})\cdot\nabla u)\right)\,\mathsf{d}\bm{\mathsf{x}}\right),$ (10)
@@
$\Re\left(\int_{\Omega}\left(\nabla u\cdot\nabla\overline{v}+2V(\bm{\mathsf{x}})u\overline{v}+2\beta|u|^{2}u\overline{v}-2i\omega\overline{v}(\bm{\mathsf{A}}(\bm{\mathsf{x}})\cdot\nabla u)\right)\,\mathsf{d}\bm{\mathsf{x}}\right)=\lambda\Re\left(\int_{\Omega}u\overline{v}\,\mathsf{d}\bm{\mathsf{x}}\right)$ (11)
for all $v\in\mathbb{H}$. Equivalently, see [31, Prop. A.1], a solution $(\lambda,u)$ of (11) satisfies
@@
meaning that it does not matter whether or not the real part is applied. This is the weak form of the stationary GPE with angular momentum, which can also be written as a semi-linear eigenfunction problem, namely, in strong form $(\lambda,u)\in\mathbb{R}\times\mathrm{C}_{0}^{2}(\Omega)$ satisfies
$-\Delta u+2V(\bm{\mathsf{x}})u+2\beta|u|^{2}u-2i\omega(\bm{\mathsf{A}}(\bm{\mathsf{x}})\cdot\nabla u)=\lambda u.$ (13)
An $\mathrm{L}^{2}(\Omega)$-normalised solution $u\in\mathbb{H}$ of (13) with minimal energy is termed a ground-state, denoted by $u_{\mathrm{GS}}\equiv u$, cf. (7), whereas any $\mathrm{L}^{2}(\Omega)$-normalised $u\in\mathbb{H}$ solving (13), with an energy strictly higher than the ground-state energy, i.e., $\mathsf{E}(u)>\mathsf{E}(u_{\mathrm{GS}})=\min_{v\in\mathcal{S}_{\mathbb{H}}}\mathsf{E}(v),$ is called an excited state.
@@
2.3. Continuous projected Sobolev gradient flow
In this section, we aim to establish a projected Sobolev gradient flow formulation for the minimisation of the Gross–Pitaevskii energy functional $\mathsf{E}$ from (1). For this purpose, we will consider the real framework of the GPE (12) rather than the complex Hilbert space approach. We emphasise, however, that both avenues ultimately result in the identical gradient flow.
@@ 2.3.1. Weighted energy inner-product
cf. [30, Eq. (7)], which will be investigated in the ensuing result.
@@
For any given $z\in\mathbb{H}$, the map $a_{z}:\mathbb{H}\times\mathbb{H}\to\mathbb{C}$ from (14) defines a (complex) inner-product on the space $\mathbb{H}=\mathrm{H}_{0}^{1}(\Omega;\mathbb{C})$. Moreover, the induced norm is equivalent to the $\mathbb{H}$-norm defined in (3); especially, for a constant $c>0$ independent of $z\in\mathbb{H}$, the uniform coercivity property $a_{z}(u,u)\geq c\left\|\nabla u\right\|_{\mathrm{L}^{2}(\Omega)}^{2}$ for all $u\in\mathbb{H}$ holds true.
@@
We will recall several ideas from [18, §2.1]. Throughout, let $z\in\mathbb{H}$ be arbitrary.
@@
    Bilinearity and symmetry. First, we note that the mapping $u\mapsto a_{z}(u,v)$ is linear for any fixed $v\in\mathbb{H}$. Moreover, we have that $a_{z}(u,v)=\overline{a_{z}(v,u)}$ for any $u,v\in\mathbb{H}$. Indeed, since $\beta\in\mathbb{R}$ and $V\in\mathrm{L}^{\infty}(\Omega;\mathbb{R})$, the conjugate-symmetry is directly observed for the first three terms in the integral from (14). Moreover, for the last term, we exploit the fact that $u$ and $v$ vanish along the boundary of $\Omega$, and hence the application of the divergence theorem implies that
    $i\omega\int_{\Omega}\overline{v}(\bm{\mathsf{A}}\cdot\nabla u)\,\mathsf{d}\bm{\mathsf{x}}=i\omega\int_{\Omega}u\operatorname*{div}(\overline{v}\bm{\mathsf{A}})\,\mathsf{d}\bm{\mathsf{x}}=i\omega\int_{\Omega}u(\nabla\overline{v}\cdot\bm{\mathsf{A}})\,\mathsf{d}\bm{\mathsf{x}}=\overline{-i\omega\int_{\Omega}\overline{u}(\bm{\mathsf{A}}\cdot\nabla v)\,\mathsf{d}\bm{\mathsf{x}}},$
@@
    $\varepsilon:=\frac{\omega^{2}r_{\Omega}^{2}}{\omega^{2}r_{\Omega}^{2}+2\delta_{\Omega}}\in(0,1),$ (17)
    with $\delta_{\Omega}>0$ from (4), it follows that
    $2V(\bm{\mathsf{x}})-\varepsilon^{-1}\omega^{2}r^{2}_{\Omega}\geq 0,\qquad\text{for (almost) all }\bm{\mathsf{x}}\in\Omega.$ (18)
    Exploiting that $\beta\geq 0$, we have
@@
    $c:=1-\varepsilon=\frac{2\delta_{\Omega}}{\omega^{2}r_{\Omega}^{2}+2\delta_{\Omega}}>0;$
@@
Remark 2.6.
The arguments in the proof of Proposition 2.3 can be applied to show that the energy functional $\mathsf{E}$ is coercive on $\mathbb{H}$. Indeed, applying the bound (16), with $\varepsilon\in(0,1)$ from (17), we notice that
$\mathsf{E}(u)$ $\geq\int_{\Omega}\frac{1}{2}\left(\left(1-\varepsilon\right)|\nabla u|^{2}+\left(V(\bm{\mathsf{x}})-\frac{1}{2}\varepsilon^{-1}\omega^{2}r^{2}_{\Omega}\right)|u|^{2}+\frac{\beta}{2}|u|^{4}\right)\,\mathsf{d}\bm{\mathsf{x}}\qquad\forall u\in\mathbb{H}.$
Recalling (18) and the fact that $\beta\geq 0$, it follows that
$\mathsf{E}(u)\geq c\left\|\nabla u\right\|^{2}_{\mathrm{L}^{2}(\Omega)}\qquad\forall u\in\mathbb{H},$ (22)
with the constant $c=\nicefrac{{(1-\varepsilon)}}{{2}}>0$.